Il Review Learning Goal: To be able to find the center of gravity, the center of mass, and the centroid of a composite body.I need part E please (previous incorrect answers for part e: 4.4, 4.93 and 4.7)
Answer #1

V z 279= (1:329= 9:634200 m₂ = 45 89.634 = 0.250 lb. V = Va-V6 = 1.983 in 3. m, - 10x /19830.011 47 lb. V. Va = 1/3 treho = {